# Diagnosing phase transitions through time scale entanglement

### v1

@@ V Hubbard dimer
In the case of the Hubbard dimer [$N=2,t^{\prime}=0$ in Eq. (IV)], all 16 eigenstates and -values can be analytically calculated [39] and tracked, providing a first good testcase for the QTTD approach. Fig. 1 shows the bond dimension of the QTT representation of one- and two-particle Green’s functions as well as the corresponding spin susceptibility $\chi_{S}$ in the $\mu/U-U$-plane. In Fig. 1(a), the bond dimension of the local one-particle Green’s function $G_{11}^{\uparrow\uparrow}(\tau)=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau)\hat{c}_{1,\uparrow}^{\dagger}\rangle$ is shown at $\beta=50,\epsilon=2\cdot 10^{-6},R=12$. It can be seen that due to the small eigenspace (only 16 eigenvalues) and the cutoff in the QTT compression the bond dimension is rather small, because only few exponentials contribute. However, a peak in the bond dimension can be seen in the vicinity of the ground state crossover, indicated by the dashed lines. Here, the singlet ground state at half filling crosses with a doublet that has one (three) electrons at small (large) values of $\mu/U$. The picture becomes clearer when analyzing the bond dimension of the local two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau_{1})\hat{c}_{1,\uparrow}^{\dagger}(\tau_{2})\hat{c}_{1,\uparrow}(\tau_{3})\hat{c}_{1,\uparrow}^{\dagger}\rangle$ in Fig. 1 (b) with $\epsilon=10^{-14}$ at $\beta=50$ and $R=6$. Here, for three imaginary times more exponentials contribute giving rise to a smoother picture with larger bond dimensions. Thus, a sharp peak can be seen as a remnant of the QPT. Increasing temperature ($\beta=30$) in Fig. 1(d) leads to the expected smearing in the vicinity of the transition. In Fig. 1(c), the value of the spin susceptibility $\chi_{S}$ (see Appendix for definition) is shown, clearly distinguishing the two “phases”. In the Supplemental Material [40], we show that this analysis also applies to other correlators (including Matsubara frequency dependence) and temperatures, and also to crossings of the first excited states. We additionally demonstrate how to proceed with low-accuracy data by applying higher QTT cutoffs.
@@ VI Four-site Hubbard ring
Next, let us consider the periodic 4-site Hubbard ring with nearest neighbor hopping only, i.e., $N=4,t^{\prime}=0$ in Eq. (IV). In Fig. 2(a), the bond dimension of the QTT representation of the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown in the $\mu/U-U$-plane. Crossings of the ground state energy are indicated by red dots. For not too low values of $U$ it can be seen that the dots lie precisely on top of the peaks of the bond dimension. This becomes also visible in Fig. 2(e), where the slice for $U=4$ is shown and the remnants of the QPT (red dotted lines) can be seen as sharp peaks in $D_{\mathrm{max}}$. For lower values of $U$ the bond dimension peaks deviate from the crossings in the eigenspectrum of the Hamiltonian (red dashed lines) and become broader. This is because for small $U$ and $\beta=50$ excited states become thermally accessible. Consequently, when decreasing $T$, the region with large bond dimensions shifts to smaller $U$’s, see Fig. 2(f), and for $T\rightarrow 0$ again alienates with the eigenvalue crossings.
@@
The spin susceptibility $\chi_{S}$ in Fig. 2(c) and the fidelity susceptibility $\chi_{F}$ (in Fig. 2(d) (see Appendix C for its definition) show a qualitatively similar behavior as the bond dimension: at large values of $U$ the ground state crossings are marked by a maximum, while for small values of $U$ both also show larger values, absent of an eigenvalue crossing. Further, when investigating entanglement in Fig. 2(b), also a similar structure is seen, where the white dashed line indicates the quantum Fisher information $F_{Q}=1$ (see Appendix for definition). Hence, the peaks in the bond dimension for low $U$ in Fig. 2(a) seem to be associated with some thermal crossover between different regimes. Further evidence of this thermal crossover of regimes seen by QTTD can be seen in Fig. 2(f), where the QTT bond dimension of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ is shown for various temperatures. The dashed lines indicate where $F_{Q}=1$ and are located at the maxima of the different curves. Interestingly, also the derivative of the quantum Fisher information $\frac{dF_{Q}}{dU}$ (and other entanglement measures - see SM [40]) as well as that of the double occupancy show a maximum in close vicinity of this point. Additionally, also the maximum and inflection point of the spin susceptibility lie in proximity to this point and show a similar temperature behavior. We refer the reader to the SM [40] for further details and figures on all of the properties discussed.
@@
Figure 2: Four-site Hubbard ring with nearest-neighbor hopping only: (a) $D_{\mathrm{max}}$ of the QTT for $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ at $\beta=50$, $\epsilon=10^{-14}$ and $R=6$ in comparison to (b) the quantum Fisher information $F_{Q}$, where the white dashed line indicates $F_{Q}=1$ , (c) spin susceptibility $\chi_{S}$ and (d) fidelity susceptibility. (e) and (f): $U=4$ and $\mu=U/2$ slice of (a), where red dashed lines in (e) indicate QPTs and dashed lines in (f) indicate $F_{Q}=1$. In (f), two additional $\beta$’s are shown.
@@ VII Four-site Hubbard ring with next-nearest neighbor hopping
Next, as an indicator of time scale entanglement in the system, the bond dimension of the QTT representation of the two-particle correlator $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ of the periodic 4-site Hubbard ring with next-nearest neighbor hopping $t^{\prime}$ at $U=10$, $\beta=50$ is studied in the $t^{\prime}-\mu/U$-plane [41]. In contrast to the previous examples, we apply larger QTT cutoffs $\epsilon$. In Fig. 3(a), $D_{\mathrm{max}}$ of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown for $\epsilon=10^{-5}$. Still, the different ground state crossings (indicated by red dots) between s(inglet), d(oublet) and q(uartet) states lie precisely on top of the peaks of $D_{\mathrm{max}}$. This becomes even better visible in Figs. 3(b)-(c), where different $\mu/U$-slices are shown for a different QTT cutoff of $\epsilon=10^{-8}$. Here, the sharp peaks of the bond dimension are again precisely located at the QPTs demonstrating how well the proposed QTTD method works by producing stable results for different QTT cutoffs ($\epsilon=10^{-5}$ in (a) vs. $\epsilon=10^{-8}$ in (b)-(c)). This shows that the time scale entanglement diagnosed by the growth of the bond dimension $D_{\mathrm{max}}$ is system-inherent and visible across all studied correlators (see SM [40] additional plots). In App. B, a short recipe on how QTT cutoffs can be efficiently used, to deal with low-accuracy data is presented and in the SM [40], we comment on crossings that are correctly diagnosed by QTTD for which $\chi_{F}$ fails to achieve the same.
@@ VIII Single-impurity Anderson model
We examine the single-impurity Anderson model (SIAM), which features a magnetic impurity in a metallic bath and shows, among others, a crossover between local moment and Kondo regime. For numerical calculations, the continuous bath is discretized leading to the following Hamiltonian
@@
Here, $\varepsilon_{0}$ represents the impurity site energy, $\mu$ is the chemical potential, and $U$ is the on-site Coulomb interaction on the impurity site. The annihilation (creation) operator on the impurity site is denoted by $\hat{d}_{\sigma}^{({\dagger})}$, and the number operator is $\hat{n}_{\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$. The density of states of the bath is assumed to be semicircular, and the parameters $V_{\ell}$ and $E_{\ell}$ denote the hybridization strength and energy levels of the discretized bath, respectively. The continuous bath is discretized into three levels, as detailed in the SM [40]. Figure 4(a) illustrates the bond dimension of the QTT representation of the two-particle Green’s function $G^{\uparrow\uparrow\uparrow\uparrow}_{\mathrm{imp}}(\tau_{1},\tau_{2},\tau_{3})$ at the impurity site for $\beta=100$, plotted against the model parameters $U$ and $V$, where $V$ denotes the hybridization strength in the original continuous model. In Fig. 4(b), the spin susceptibility $\chi_{S}$ is depicted. The red dots mark the Kondo temperature $T_{K}=V\sqrt{U}e^{-\pi U/8V^{2}}$ estimated via poor man’s scaling [42, 43], which delineates the boundary between the Kondo and local moment regimes, as also reflected in the spin susceptibility behavior. The bond dimension has a broad peak located in the vicinity of the Kondo temperature, diagnosing the thermally driven Kondo to local moments regime crossover.
Figure 4: (a) Bond dimension $D_{\mathrm{max}}$ of the two-particle Green’s function of the single-impurity Anderson model at $\beta=100$ in the $U$-$V$ plane, where the red dots indicate $T_{K}(U,V)=1/\beta$. (b) Spin susceptibility $\chi_{S}$ indicating the different regimes.
@@ IX Conclusions
Since their recent invention, QTTs have primarily been applied as a compression tool in numerical calculations to overcome severe memory bottlenecks. In this work, we have presented an entirely novel approach on diagnosing phase transitions and crossovers with QTTs. We have demonstrated that this new quantics tensor train diagnostics approach is able to identify QPTs and thermal regime crossovers and distinguish between these two by studying various one-dimensional Hubbard rings and the SIAM. Moreover, by showing results for different correlators we have identified time scale entanglement to be system-inherent showing maxima at phase transitions and crossovers and, additionally, have pointed out that this method is also able to deal with low-accuracy data by varying the QTT cutoff. For the four-site ring with next-nearest neighbor hopping, QTTD correctly diagnosed ground state crossings that the fidelity susceptibility failed to capture. The strong point of the QTTD is that it can be used universally for zero and finite temperature and multiple correlation functions – it is not limited to ground state methods and does not require specific measures or susceptibilities. Further, QTTD offers a new perspective and deeper understanding on the inherent time and length scale entanglement at phase transitions.
@@ X Acknowledgements
We thank Samuel Badr, Jan von Delft, Herbert Eßl, Markus Frankenbach, Matthias Reitner, Marc Ritter and Nepomuk Ritz for insightful discussions. This work was funded in part by the Austrian Science Fund (FWF) projects through Grant DOI 10.55776/P36332, 10.55776/F86, 10.55776/I5868, and 10.55776/V1018. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. Calculations have been partly performed on the Vienna Scientific Cluster (VSC). H.S. was supported by JSPS KAKENHI Grants No. 21H01041, No. 21H01003, and No. 23H03817 as well as JST FOREST Grant No. JPMJFR2232, Japan. H.I. and H.S. were supported by JST PRESTO Grant No. JPMJPR2012, Japan.
@@ Appendix A Appendix A: QTTD and multi-point correlators
Let us consider an arbitrary n-point correlator function $\mathcal{G}(\bm{\tau})=-\langle\mathcal{T}\hat{O}_{1}(\tau_{1})\hat{O}_{2}(\tau_{2})...\hat{O}_{n}(0)\rangle$ with $\bm{\tau}=(\tau_{1},\tau_{2},...,\tau_{n-1}),\tau_{i}\in[0,\beta)$ and fermionic or bosonic operators $\hat{O}_{i}$. We express the correlator in terms of its Lehmann representation
$\mathcal{G}(\bm{\tau})$ $=-\frac{1}{Z}\sum_{m_{1}}e^{-\beta\epsilon_{m_{1}}}\sum_{m_{2}...m_{n}}e^{-\tau_{1}(\epsilon_{m_{2}}-\epsilon_{m_{1}})}e^{-\tau_{2}(\epsilon_{m_{3}}-\epsilon_{m_{2}})}...e^{-\tau_{n-1}(\epsilon_{m_{n}}-\epsilon_{m_{n-1}})}$
×$\times\langle m_{1}|\hat{O}_{1}|m_{2}\rangle\langle m_{2}|\hat{O}_{2}|m_{3}\rangle...\langle,m_{n}|\hat{O}_{n}|m_{1}\rangle$ (7)
@@
where we already set $\epsilon_{\mathrm{GS}}=0$. Similar to the analysis of Eq. (III), the first summand dominates in the case of small $\tau_{1},\tau_{2}$, the second one in the case of large $\tau_{1}$ (close to $\beta$) and small $\tau_{2}$ and the third for large $\tau_{1},\tau_{2}$. In contrast to the case of the two-point correlator, it can be seen that apart from the dependence on two instead of one imaginary times, in general, more exponentials will contribute in the three-point correlator. Since the QTT representation of an exponential has a bond dimension of one and the bond dimension of the sum of two QTTs is bound by the sum of the bond dimensions of the individual QTTs, we expect three-point correlators to have larger QTT bond dimensions than two-point correlators due to additional imaginary time dependence and more dominating contributions in the sum. Therefore, higher-point correlators are also expected to show more pronounced peaks at phase transitions and crossovers, which is confirmed by the results in the paper. The reasoning that phase transitions and crossovers are connected to maxima in the bond dimensions of correlators follows the same arguments as in the main text. Let us emphasize that this derivation is valid for arbitrary (non-zero) correlators. Since Fourier transformation can be represented by a low-rank matrix product operator applied to the correlator QTT, imaginary frequency counterparts to the imaginary times correlator are expected to display the same maxima in the bond dimension, which is supported by the results in the Supplemental Material. Therefore, arbitrary correlators that are available from computations can be used in QTTD to diagnose phase transitions and crossovers in systems. Furthermore, since the bond dimension is a measure of time scale entanglement, we argue that time scale entanglement is primarily a system-inherent property, because arbitrary correlators share the same bond dimension features.

### v2

@@ V Hubbard dimer
For Hubbard dimer $N=2,t^{\prime}=0$ in Eq. (5)], all 16 eigenstates can be analytically calculated [48], providing a first good testcase for QTTD. Fig. 1(b) shows $D_{\mathrm{max}}$ of the local two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau_{1})\hat{c}_{1,\uparrow}^{\dagger}(\tau_{2})\hat{c}_{1,\uparrow}(\tau_{3})\hat{c}_{1,\uparrow}^{\dagger}\rangle$ with $\epsilon=10^{-14}$ at $\beta=50$ and $R=6$. A sharp peak in $D_{\mathrm{max}}$ can be seen in the vicinity of the ground state crossover, indicated by dashed lines. Here, the singlet ground state crosses with a doublet that has one (three) electrons at small (large) values of $\mu/U$. In Fig. 1(a), the value of the spin susceptibility $\chi_{S}$ (definition in App. E) is shown, clearly distinguishing the two “phases”. Fig. 1(c) plots the bond dimension as time scale entanglement measure over the imaginary time scale. In App. A and the Supplemental Material [49], we show that this analysis also applies to other correlators (including Matsubara frequency dependence) and temperatures, and also to crossings of the first excited states. We additionally demonstrate how to proceed with low-accuracy data by applying higher QTT cutoffs.
@@ VI Four-site Hubbard ring
We consider the periodic 4-site Hubbard ring with nearest neighbor hopping only, i.e., $N=4,t^{\prime}=0$, $t_{\pm(N-1)}=t$, in Eq. (5). In Fig. 2(a), $D_{\mathrm{max}}$ of the QTT representation of the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown in the $\mu/U-U$-plane. Crossings of the ground states are indicated by red dots. For not too low values of $U$ the dots lie precisely on top of the peaks of $D_{\mathrm{max}}$. In Fig. 2(e), the slice for $U=4$ is shown. For lower values of $U$ the $D_{\mathrm{max}}$ peaks deviate from the crossings in the eigenspectrum of the Hamiltonian (red dashed lines) and become broader. This is because for small $U$ and $\beta=50$ excited states become thermally accessible. Consequently, when decreasing $T$, the region with large $D_{\mathrm{max}}$ shifts to smaller $U$’s, see Fig. 2(f), and for $T\rightarrow 0$ again alienates with the eigenvalue crossings.
@@
The spin susceptibility $\chi_{S}$ in Fig. 2(c) and, e.g., the fidelity susceptibility $\chi_{F}$ (in Fig. 2(d) (definition in App. C) show a qualitatively similar behavior as $D_{\mathrm{max}}$: at large values of $U$ the ground state crossings are marked by a maximum, while for small values of $U$ both show larger values, without an eigenvalue crossing. Further, when investigating entanglement in Fig. 2(b), also a similar structure is seen, where the white dashed line indicates the quantum Fisher information $F_{Q}=1$ (definition in App. F). Hence, the peaks in $D_{\mathrm{max}}$ for low $U$ in Fig. 2(a) seem to be associated with some thermal crossover between different regimes. Further evidence of this thermal crossover seen by QTTD is seen in Fig. 2(f), where the QTT bond dimension of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ is shown for various temperatures. The dashed lines indicate $F_{Q}=1$ and are located at the maxima of the different curves. Interestingly, also the derivative $\frac{dF_{Q}}{dU}$ (and other entanglement measures - see SM [49]) as well as that of the double occupancy show a maximum in close vicinity of this point. Additionally, also the maximum and inflection point of the spin susceptibility lie in proximity to this point and show a similar temperature behavior (further details in SM [49]).
@@
Figure 2: Four-site Hubbard ring with nearest-neighbor hopping only: (a) $D_{\mathrm{max}}$ of the QTT for $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ at $1/T=\beta=50$, $\epsilon=10^{-14}$ and $R=6$ in comparison to (b) the quantum Fisher information $F_{Q}$, where the white dashed line indicates $F_{Q}=1$ , (c) spin susceptibility $\chi_{S}$ and (d) fidelity susceptibility. (e) and (f): $U=4$ and $\mu=U/2$ slice of (a), where red dashed lines in (e) indicate QPTs and dashed lines in (f) indicate $F_{Q}=1$. In (f), two additional $\beta$’s are shown.
@@ VII Four-site Hubbard ring with next-nearest neighbor hopping
We consider $D_{\mathrm{max}}$ of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ of the periodic 4-site Hubbard ring with next-nearest neighbor hopping $t^{\prime}$ at $U=10$, $\beta=50$ in the $t^{\prime}-\mu/U$-plane [50]. Contrary to the previous examples, we apply larger QTT cutoffs $\epsilon$. In Fig. 3(a), $D_{\mathrm{max}}$ of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown for $\epsilon=10^{-5}$. Still, the different ground state crossings (indicated by red dots) between s(inglet), d(oublet) and q(uartet) states lie precisely on top of the peaks of $D_{\mathrm{max}}$.
In Figs. 3(b)-(c), different $\mu/U$-slices are shown for a lower QTT cutoff of $\epsilon=10^{-8}$, demonstrating how well the proposed QTTD works by producing stable $D_{\mathrm{max}}$ peaks for different QTT cutoffs ($\epsilon=10^{-5}$ in (a) vs. $\epsilon=10^{-8}$ in (b)-(c)). This shows that time scale entanglement diagnosed by the growth of $D_{\mathrm{max}}$ is system-inherent and visible across all studied correlators (see SM [49] additional plots). In the SM [49], we comment on crossings that are correctly diagnosed by QTTD for which $\chi_{F}$ fails to achieve the same.
@@ VIII Single-impurity Anderson model
The single-impurity Anderson model (SIAM) features a magnetic impurity in a metallic bath and shows a crossover between local moment and Kondo regime. For numerical calculations, the continuous bath is discretized leading to the following Hamiltonian
@@
$\varepsilon_{0}$ represents the impurity site energy, $\mu$ is the chemical potential, and $U$ the on-site interaction. The annihilation (creation) operator on the impurity site is $\hat{d}_{\sigma}^{({\dagger})}$, and $\hat{n}_{\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$. The bath density of states is assumed to be semicircular, and the parameters $V_{\ell}$ and $E_{\ell}$ denote the hybridization strength and energy levels of the discretized bath, respectively. The bath is discretized into three levels, as detailed in the SM [49]. Figure 4(a) illustrates $D_{\mathrm{max}}$ of the QTT representation of $G^{\uparrow\uparrow\uparrow\uparrow}_{\mathrm{imp}}(\tau_{1},\tau_{2},\tau_{3})$ at the impurity site for $\beta=100$, plotted against the model parameters $U$ and $V$, where $V$ denotes the hybridization strength in the original continuous model. In Fig. 4(b), the spin susceptibility $\chi_{S}$ is depicted. The red dots mark the Kondo temperature $T_{K}=V\sqrt{U}e^{-\pi U/8V^{2}}$ estimated via poor man’s scaling [51, 52], which delineates the boundary between Kondo and local moment regimes, as also reflected in the spin susceptibility behavior. $D_{\mathrm{max}}$ has a broad peak located in the vicinity of the Kondo temperature, diagnosing the thermally driven Kondo to local moments regime crossover.
Figure 4: (a) $D_{\mathrm{max}}$ of the two-particle Green’s function of the SIAM at $\beta=100$ in the $U$-$V$ plane [red dots indicate $T_{K}(U,V)=1/\beta$]. (b) Spin susceptibility $\chi_{S}$ indicating the different regimes.
@@ IX Conclusions
Since their recent invention, QTTs have primarily been applied as a compression tool in numerical calculations to overcome severe memory bottlenecks. In this work, we present a novel QTTD approach, diagnosing time-scale entanglement from the QTT bond dimension. We further demonstrate that QTTD is actually able to identify thermal crossovers, ground state changes, and presumably QPTs, and to distinguish between these phenomena.
Studying various one-dimensional Hubbard rings and the SIAM, we show that time-scale entanglement for all correlators becomes maximal at QPTs and crossovers. This is even true for the one-particle Green’s function whose static, time-independent part does not signal a second order phase transition when approaching it from the symmetry-unbroken side. In the Green’s function dynamics, the phase transition is however encoded in its time-scale entanglement. Additionally, we point out that QTTD is able to deal with low-accuracy data by varying the QTT cutoff. Comparing to conventional entanglement measures, we show that QTTD correctly diagnoses ground state crossings that the fidelity susceptibility fails to capture.
The strong point of QTTD is that it can be used universally for zero and finite temperature and multiple correlation functions—it is not limited to ground state methods and does not require specific measures or susceptibilities. Further, QTTD offers a new perspective and deeper understanding on the inherent time and length scale entanglement at phase transitions. It can provide such a fresh perspective also for other physical phenomena.
@@ X Acknowledgements
We thank Samuel Badr, Jan von Delft, Herbert Eßl, Markus Frankenbach, Matthias Reitner, Marc Ritter and Nepomuk Ritz for insightful discussions. This work was funded in part by the Austrian Science Fund (FWF) projects through Grant DOI 10.55776/P36332, 10.55776/F86, 10.55776/I5868, and 10.55776/V1018. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. Calculations have been partly performed using Austrian Scientific Computing (ASC). H.S. was supported by JSPS KAKENHI Grants No. 21H01041, No. 21H01003, and No. 23H03817 as well as JST FOREST Grant No. JPMJFR2232, Japan.
@@ Appendix A Appendix A: QTTD and multi-point correlators
Let us consider an arbitrary n-point correlator function $\mathcal{G}(\bm{\tau})=-\langle\mathcal{T}\hat{O}_{1}(\tau_{1})\hat{O}_{2}(\tau_{2})...\hat{O}_{n}(0)\rangle$ with $\bm{\tau}=(\tau_{1},\tau_{2},...,\tau_{n-1}),\tau_{i}\in[0,\beta)$ and fermionic or bosonic operators $\hat{O}_{i}$. For $\tau_{1}>\tau_{2}>\ldots\tau_{n-1}$, we express the correlator in terms of its Lehmann representation:
$\mathcal{G}(\bm{\tau})$ $=-\frac{1}{Z}\sum_{m_{1}\ldots m_{n}}e^{-\tau_{1}(\epsilon_{m_{2}}-\epsilon_{m_{1}})}e^{-\tau_{2}(\epsilon_{m_{3}}-\epsilon_{m_{2}})}\cdots e^{-\tau_{n-1}(\epsilon_{m_{n}}-\epsilon_{m_{n-1}})}$
$\times e^{-\beta\epsilon_{m_{1}}}\langle m_{1}|\hat{O}_{1}|m_{2}\rangle\langle m_{2}|\hat{O}_{2}|m_{3}\rangle\cdots\langle m_{n}|\hat{O}_{n}|m_{1}\rangle$ (7)
@@
where we already set $\epsilon_{\mathrm{GS}}=0$. Similar to the analysis of Eq. (III), the first summand dominates in the case of small $\tau_{1},\tau_{2}$, the second one in the case of large $\tau_{1}$ (close to $\beta$) and small $\tau_{2}$ and the third for large $\tau_{1},\tau_{2}$. In contrast to the case of the two-point correlator, it can be seen that apart from the dependence on two instead of one imaginary times, in general, more exponentials will contribute in the three-point correlator. Since the QTT representation of an exponential has a bond dimension of one and the bond dimension of the sum of two QTTs is bound by the sum of the bond dimensions of the individual QTTs, we expect three-point correlators to have larger QTT bond dimensions than two-point correlators due to additional imaginary time dependence and more dominating contributions in the sum. Therefore, higher-point correlators are also expected to show more pronounced peaks at phase transitions and crossovers (cf. Fig. 1 and Fig. 5). The reasoning that phase transitions and crossovers are connected to maxima in the bond dimensions of correlators follows the same arguments as in the main text. Let us emphasize that this derivation is valid for arbitrary (non-zero) correlators. Since Fourier transformation can be represented by a low-rank matrix product operator [38, 53] applied to the correlator QTT, imaginary frequency counterparts to the imaginary times correlator are expected to display the same maxima in the bond dimension, which is supported by the results in the Supplemental Material. Therefore, arbitrary correlators that are available from computations can be used in QTTD to diagnose phase transitions and crossovers in systems. Furthermore, since the bond dimension is a measure of time scale entanglement, we argue that time scale entanglement is primarily a system-inherent property, because arbitrary correlators share the same bond dimension features.